OpenStudy (anonymous):

determine whether the lines through each pair are perpendicular? (10,6); (-4,-3); and (1,-7), I said it's not, is this correct?

myininaya (myininaya):

so (10,6) amd (-4,-3) is on one line? and so (1,-7) and ______ is on the other line?

OpenStudy (anonymous):

Sorry, I missted something? it's (10,6) and (2,16) (-4,-3) and (1,-7)

myininaya (myininaya):

slope of line 1 m1=(16-6)/(2-10)=10/(-8)=-10/8=-5/4 slope of line 2 m2=(-7-(-3))/(1-(-4))=-4/5 so no they aren't both slopes are negative so not perpendicular